OpenStudy (anonymous):

if 13824=6^x times 4^y what is the value of xtimes y?

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

(6^3)(4^3)=216*64=13824 x=3 y=3 x*y=3*3=9

OpenStudy (anonymous):

how do u get that?

OpenStudy (anonymous):

i know that anything divisible by 6 has to b divisible by 3 because 3*2=6 so divide it as much times as u can by 3 and u kno how many 6's there are. ( u can also divide by 6 as much times as u can u get the same answer). after that u kno whatever is left over must b a power of 4